OpenStudy (anonymous):

I need help!! Find a quadratic equation with roots -1+4i and -1-4i

OpenStudy (prettygirl_shynice):

0k

OpenStudy (xapproachesinfinity):

if a and b are root of a quadratic then (x-a)(x-b)=0

OpenStudy (xapproachesinfinity):

so you are given two roots what do you do?

OpenStudy (anonymous):

plug the values in and distribute?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

Thank you sooo much!!!! :D

OpenStudy (solomonzelman):

This is how I would prove. (b is a positive constant) \(\large\color{#000000 }{ \displaystyle x=-a\pm bi }\) \(\large\color{#000000 }{ \displaystyle x+a=\pm bi }\) \(\large\color{#000000 }{ \displaystyle (x+a)^2=-b }\) \(\large\color{#000000 }{ \displaystyle x^2+2ax-a^2=-b }\) \(\large\color{#000000 }{ \displaystyle x^2+2ax-a^2+b =0}\)

OpenStudy (solomonzelman):

completing the square, but, backwards ))